OpenStudy (anonymous):

-1/9 + (-5/7) (-1/9 and -5/7 are decimals)

OpenStudy (anonymous):

actually they are fractions

OpenStudy (anonymous):

\[-\frac{1}{9}-\frac{5}{7}=\frac{-7-5\times 9}{9\times 7}\]

OpenStudy (anonymous):

oops i mean fractions